# Learning Dense Hand Contact Estimation from Imbalanced Data

### v1

@@ Appendix A Appendix
  • A.6. Quantitative results with different backbones
  • A.7. Computational requirements
  • A.8. More qualitative results
  • A.9. Limitations and societal impacts
@@ A.1 Details of loss functions
Smoothness loss. We penalize spatially isolated or fragmented contact predictions by leveraging the mesh topology of the hand. Let $p_{v}\in[0,1]$ denote the predicted contact probability after the sigmoid for each vertex $v\in V$, and let $\mathbf{A}\in\{0,1\}^{V\times V}$ be a sparse adjacency matrix representing vertex connectivity in the mesh. Here, $A_{vu}=1$ indicates that vertex $u$ is connected to vertex $v$, and 00 otherwise. For each vertex $v\in V$, we compute the aggregated contact prediction over its neighbors as:
@@
where $\hat{p}_{v}$ and $\hat{q}_{v}$ represent the neighborhood-smoothed estimates for contact and non-contact probabilities, respectively. To assess spatial inconsistency, we compute the discrepancy between each vertex’s prediction and its neighborhood average. Specifically, we define the isolation score as:
@@
which corresponds to the degree of vertex $v$ in the mesh. This normalization accounts for how many vertices contribute to the isolation score at each location, ensuring that the loss measures average inconsistency rather than accumulating errors over larger neighborhoods. As a result, the model is encouraged to produce smooth contact maps without being biased against large, spatially coherent contact regions. We then define the final smoothness regularization loss as:
@@ A.2 Details of contact initialization
To stabilize the training of dense hand contact estimation of HACO, we propose a contact initialization technique that utilizes a learnable embedding to learn the most effective initial contact for dense hand contact estimation. In Table A1, our proposed contact initialization outperforms all other initialization methods, achieving the superior performance in precision, recall, and F1-score. Unlike other approaches that rely on fixed priors or simplistic assumptions, our method introduces a contact representation learned from input data, enabling more accurate and robust initialization during inference. This formulation captures the most useful initial prediction compared to static or handcrafted methods, leading to gains across all metrics. Such consistent improvement demonstrates that learning contact directly from large-scale datasets, rather than relying on pre-defined numbers or dataset averages, is crucial for precise contact estimation.
Table A1: Comparison of various contact initialization on MOW cao2021reconstructing dataset.
[…]
@@
Mean of ObMan hasson2019learning 0.520 0.610 0.522
Mean of DexYCB chao2021dexycb 0.520 0.588 0.512
Mean of MOW cao2021reconstructing 0.511 0.600 0.515
Mean of IH26M moon2020interhand2 0.522 0.567 0.505
Ours 0.525 0.632 0.531
@@ A.3 Details of dense hand contact labels
Following the previous works on dense human-scene contact estimation hassan2021populating ; huang2022capturing ; nam2024joint , we employ distance-based thresholding to generate ground-truth dense hand contact labels. This is implemented using the Trimesh library trimesh . Given a ground-truth hand mesh and the mesh of the interacting entity (i.e., an object for hand-object interaction, another hand for hand-hand interaction, a face for hand-face interaction, the environment for hand-scene interaction, or the body for hand-body interaction), we construct a proximity query from the interacting mesh and compute the closest surface point on it for each vertex of the hand mesh. Contact is then determined by thresholding the Euclidean distance between each hand vertex and its corresponding closest point. The resulting contact labels are binary values per vertex, indicating whether each vertex is in contact.
To set these thresholds, we manually inspected the ground-truth meshes and the resulting contact labels produced under fixed threshold values. We observed that the accuracy of 3D annotations varies significantly across datasets, making a single threshold unsuitable for all cases. A uniform threshold often led to incorrect labeling and semantically implausible contact regions. Based on our analysis, we set the contact distance threshold to 1 cm for the following datasets: ObMan hasson2019learning , DexYCB chao2021dexycb , HO3D hampali2020honnotate , H2O3D hampali2022keypoint , ARCTIC fan2023arctic , HOI4D liu2022hoi4d , H2O kwon2021h2o , and PROX hassan2019resolving . For the MOW cao2021reconstructing dataset, we used a threshold of 3.5 cm due to its lower mesh fidelity. For InterHand2.6M moon2020interhand2 and HIC tzionas2016capturing , we adopted a smaller threshold of 0.5 cm to better capture fine-grained hand-hand interactions.
Although most datasets produced reasonable contact labels with a 1 cm threshold, MOW required a larger value of 3.5 cm due to coarse mesh annotations generated through optimization on in-the-wild images. For hand-hand interaction datasets, the 1 cm threshold, which is approximately the width of a finger, often resulted in inaccurate and semantically implausible contact labels. To address this, we adopted a smaller threshold of 0.5 cm. For datasets with existing annotated contact labels, such as RICH huang2022capturing , Decaf shimada2023decaf , and Hi4D yin2023hi4d , we directly used the provided ground-truth annotations.
Figure A1: More examples of class imbalance between contact and non-contact on H2O kwon2021h2o , H2O3D hampali2022keypoint , HIC tzionas2016capturing , MOW cao2021reconstructing , ObMan hasson2019learning , PROX hassan2019resolving , RICH huang2022capturing , Hi4D yin2023hi4d .
@@ A.4 More examples of class imbalance issue
Figure A1 presents the class imbalance issue across datasets kwon2021h2o ; hampali2022keypoint ; tzionas2016capturing ; cao2021reconstructing ; hasson2019learning ; hassan2019resolving ; huang2022capturing ; yin2023hi4d omitted from the main paper due to space constraints. In general, all datasets exhibit a severe class imbalance, with the majority of samples corresponding to non-contact regions. Notably, some datasets such as H2O and HIC display highly skewed ratios of 10.57:1 and 58.41:1, respectively, between non-contact and contact samples. It is important to note that contact and non-contact counts are computed at the vertex level rather than per hand instance. While addressing class imbalance may appear straightforward, the sampling process must not only balance class proportions but also preserve the diversity of contact configurations present in the dataset. To tackle this challenge, we introduce Balanced Contact Sampling (BCS), which constructs multiple sampling groups that fairly represent both contact and non-contact. Please refer to the main paper for further details on the BCS.
@@ A.5 More examples of spatial imbalance issue
We visualize the dataset-wide mean contact maps for all dense hand contact datasets used in training HACO in Figure A2 and Figure A3. Among hand-object interaction datasets, DexYCB chao2021dexycb , HO3D hampali2020honnotate , H2O3D hampali2022keypoint , ARCTIC fan2023arctic , HOI4D liu2022hoi4d , and H2O kwon2021h2o exhibit significant spatial imbalance, with ground-truth contact concentrated predominantly on the fingertips. ObMan hasson2019learning and MOW cao2021reconstructing display more promising contact distributions, including high contact probability on the hypothenar region. However, ObMan is a synthetic dataset with domain gap towards real images, and MOW, while showing diverse palmar contact patterns, lacks dorsal contact and contains fewer than 1,000 samples, making it insufficient for large-scale training. Among hand-hand interaction datasets, HIC tzionas2016capturing shows severe spatial imbalance, with contact limited to small fingertip regions and moderate values between fingers. In contrast, InterHand2.6M moon2020interhand2 offers a more favorable contact distribution, covering nearly all hand regions except for a small area on the dorsum. Hand-scene interaction datasets hassan2019resolving ; huang2022capturing tend to involve large contact areas, primarily on the palmar side. However, interactions are often limited to flat surfaces such as the ground or walls, resulting in low diversity of contact types despite broad coverage. The hand-face interaction dataset Decaf shimada2023decaf also exhibits poor spatial diversity due to its limited action types, such as poking or touching the chin. For hand-body interactions, Hi4D yin2023hi4d includes extensive contact regions, but most interactions involve full palmar contact during actions such as hugging or patting. As visualized in the dataset-wide mean contact heatmaps, spatial imbalance is a prevalent issue across existing dense hand contact datasets. To address this, we propose the vertex-level class-balanced (VCB) loss, which mitigates spatial imbalance by reweighting the loss at each vertex based on its mean contact frequency across the dataset. Please refer to the main paper for further details on the VCB loss.
Table A2: Comparison of various backbone models on MOW cao2021reconstructing dataset.
[…]
HACO ViT-H dosovitskiy2020image HaMeR pavlakos2024reconstructing 0.525 0.632 0.531
HACO FPN lin2017feature HandOccNet park2022handoccnet 0.489 0.551 0.473
HACO ViT-L dosovitskiy2020image ImageNet deng2009imagenet 0.513 0.616 0.519
HACO ViT-B dosovitskiy2020image ImageNet deng2009imagenet 0.498 0.666 0.526
HACO ViT-S dosovitskiy2020image ImageNet deng2009imagenet 0.519 0.571 0.495
HACO HRNet-W48 sun2019deep ImageNet deng2009imagenet 0.494 0.587 0.491
HACO HRNet-W32 sun2019deep ImageNet deng2009imagenet 0.492 0.594 0.486
HACO ResNet-152 he2016deep ImageNet deng2009imagenet 0.516 0.604 0.509
HACO ResNet-101 he2016deep ImageNet deng2009imagenet 0.503 0.589 0.499
HACO ResNet-50 he2016deep ImageNet deng2009imagenet 0.508 0.596 0.493
HACO ResNet-34 he2016deep ImageNet deng2009imagenet 0.494 0.616 0.493
HACO ResNet-18 he2016deep ImageNet deng2009imagenet 0.496 0.598 0.485
@@
Figure A2: Dataset-wide dense hand contact mean. These heatmaps show average hand contact of ground-truths from ObMan hasson2019learning , DexYCB chao2021dexycb , MOW cao2021reconstructing , HO3D hampali2020honnotate , H2O3D hampali2022keypoint , ARCTIC fan2023arctic , HOI4D liu2022hoi4d dataset.
@@
Figure A3: Dataset-wide dense hand contact mean. These heatmaps show average hand contact of ground-truths from H2O kwon2021h2o , InterHand2.6M moon2020interhand2 , HIC tzionas2016capturing , PROX hassan2019resolving , RICH huang2022capturing , Decaf shimada2023decaf , Hi4D yin2023hi4d dataset.
A.6 Quantitative results with different backbones
@@
We evaluate the performance of HACO with various backbone architectures, keeping the rest of the pipeline fixed. As shown in Table A2, our original model using ViT-H dosovitskiy2020image pretrained on HaMeR pavlakos2024reconstructing achieves the highest F1-score, demonstrating the strongest overall performance. ViT-based backbones consistently outperform convolutional alternatives, with ViT-B achieving the highest recall and ViT-S maintaining strong precision. ViT-L also performs competitively, achieving a strong balance between precision and recall. This highlights the effectiveness of Transformer vaswani2017attention -based architectures for dense contact estimation, likely due to their global receptive fields and capacity to model long-range dependencies. To assess the benefit of pretraining on related tasks, we additionally train and evaluate a variant with FPN lin2017feature backbone pretrained with HandOccNet park2022handoccnet , a 3D hand mesh reconstruction model. This configuration yields the lowest F1-score among all backbones, suggesting that reconstruction-specific pretraining alone is insufficient for accurate hand contact prediction. HRNet models, widely used in 3D human mesh reconstruction lin2021end ; lin2021mesh ; li2022cliff ; zhang2023pymaf ; li2025hybrik , achieve moderate performance in our setting, with F1-scores below 0.5 despite their high-resolution design. ResNet backbones he2016deep exhibit a wide performance range, with deeper variants such as ResNet-152 performing relatively well, while lighter variants such as ResNet-18 offer improved efficiency at the cost of reduced accuracy. Overall, these results indicate that the performance gains of our final HACO model arise from the combination of the strong representational capacity of the ViT backbone and task-specific knowledge transferred from 3D hand mesh reconstruction. All model variants will be publicly released.
A.7 Computational requirements

### v2

@@ Appendix A Appendix
  • A.6. More comparisons with state-of-the-art methods
  • A.7. More ablation studies
  • A.8. Quantitative results with different backbones
  • A.9. Computational requirements
  • A.10. More qualitative results
  • A.11. Limitations and societal impacts
@@ A.1 Details of loss functions
Smoothness loss. We penalize spatially isolated or fragmented contact predictions by leveraging the mesh topology of the hand. Let $p_{v}\in[0,1]$ denote the predicted contact probability after the sigmoid for each vertex $v\in V$, and let $\mathbf{A}\in\{0,1\}^{V\times V}$ be a sparse adjacency matrix representing vertex connectivity in the mesh. Here, $A_{vu}=1$ indicates that vertex $u$ is connected to vertex $v$, and 0 otherwise. For numerical stability, we also include self-connections so that each vertex is connected to itself. For each vertex $v\in V$, we compute the aggregated contact prediction over its neighbors as:
@@
where $\hat{p}_{v}$ and $\hat{q}_{v}$ represent the neighborhood-smoothed estimates for contact and non-contact probabilities, respectively. To maintain consistent scale across batches, we further normalize the aggregated contact and non-contact maps by their maximum values within each batch before computing the isolation score. To assess spatial inconsistency, we compute the discrepancy between each vertex’s prediction and its neighborhood average. Specifically, we define the isolation score as:
@@
which corresponds to the degree of vertex $v$ in the mesh. This normalization accounts for how many vertices contribute to the isolation score at each location, ensuring that the loss measures average inconsistency rather than accumulating errors over larger neighborhoods. As a result, the model is encouraged to produce smooth contact maps without being biased against large spatially coherent contact regions. We then define the final smoothness regularization loss as:
@@ A.2 Details of contact initialization
To stabilize the training of HACO on dense hand contact estimation, we propose a contact initialization technique that utilizes a learnable embedding to learn the most effective initial contact for dense hand contact estimation. In Table A1, our proposed contact initialization outperforms all other initialization methods, achieving the superior performance in precision, recall, and F1-score. Unlike other approaches that rely on fixed priors or simplistic assumptions, our method introduces a contact representation learned from input data, enabling more accurate and robust initialization during inference. This formulation captures the most useful initial prediction compared to static or handcrafted methods, leading to gains across all metrics. Such consistent improvement demonstrates that learning contact directly from large-scale datasets, rather than relying on pre-defined numbers or dataset averages, is crucial for precise contact estimation.
Table A1: Comparison of various contact initialization on MOW [4] dataset.
Methods $\text{Precision}{\uparrow}$ $\text{Recall}{\uparrow}$ $\text{F1-Score}{\uparrow}$
@@
Mean of DexYCB [5] 0.520 0.588 0.512
Mean of MOW [4] 0.511 0.600 0.515
Mean of IH26M [41] 0.522 0.567 0.505
Ours 0.525 0.607 0.522
@@ A.3 Details of dense hand contact labels
Following the previous works on dense human-scene contact estimation [18, 22, 43], we employ distance-based thresholding to generate ground-truth dense hand contact labels. This is implemented using the Trimesh library [7]. Given a ground-truth hand mesh and the mesh of the interacting entity (i.e., an object for hand-object interaction, another hand for hand-hand interaction, a face for hand-face interaction, the environment for hand-scene interaction, or the body for hand-body interaction), we construct a proximity query from the interacting mesh and compute the closest surface point on it for each vertex of the hand mesh. Contact is then determined by thresholding the Euclidean distance between each hand vertex and its corresponding closest point. The resulting contact labels are binary values per vertex, indicating whether each vertex is in contact.
To set these thresholds, we manually inspected the ground-truth meshes and the resulting contact labels produced under fixed threshold values. We observed that the accuracy of 3D annotations varies significantly across datasets, making a single threshold unsuitable for all cases. A uniform threshold often led to incorrect labeling and semantically implausible contact regions. Based on our analysis, we set the contact distance threshold to 1 cm for the following datasets: ObMan [19], DexYCB [5], HO3D [15], H2O3D [16], ARCTIC [12], HOI4D [36], H2O [24], and PROX [17]. For the MOW [4] dataset, we used a threshold of 3.5 cm due to its lower mesh fidelity. For InterHand2.6M [41] and HIC [59], we adopted a smaller threshold of 0.5 cm to better capture fine-grained hand-hand interactions.
Although most datasets produced reasonable contact labels with a 1 cm threshold, MOW required a larger value of 3.5 cm due to coarse mesh annotations generated through optimization on in-the-wild images. For hand-hand interaction datasets, the 1 cm threshold, which is approximately the width of a finger, often resulted in inaccurate and semantically implausible contact labels. To address this, we adopted a smaller threshold of 0.5 cm. For datasets with existing annotated contact labels, such as RICH [22], Decaf [55], and Hi4D [65], we directly used the provided ground-truth annotations.
Figure A1: More examples of class imbalance between contact and non-contact on H2O [24], H2O3D [16], HIC [59], MOW [4], ObMan [19], PROX [17], RICH [22], Hi4D [65].
@@ A.4 More examples of class imbalance issue
Figure A1 presents the class imbalance issue across datasets [24, 16, 59, 4, 19, 17, 22, 65] omitted from the main paper due to space constraints. In general, all datasets exhibit a severe class imbalance, with the majority of vertices corresponding to non-contact. Notably, some datasets such as H2O and HIC display highly skewed ratios of 10.57:1 and 58.41:1, respectively, between non-contact and contact vertices. It is important to note that contact and non-contact counts are computed at the vertex-level rather than per-hand instance. Since the ground-truth contact labels are provided per-hand, rather than per-vertex for each image, sampling must be performed at the hand-level. This makes it challenging to tackle class imbalance issue in per-hand manner as the class imbalance issue is present in per-vertex. Hence, it is essential to design a principled, hand-level proxy that reflects the underlying per-vertex imbalance. To tackle this challenge, we introduce balanced contact sampling (BCS), which constructs multiple sampling groups in hand-level but still fairly represent both contact and non-contact in per-vertex level. Please refer to the main paper for further details on the BCS.
@@ A.5 More examples of spatial imbalance issue
We visualize the dataset-wide mean contact maps for all dense hand contact datasets used in training HACO in Figure A2 and Figure A3. Among hand-object interaction datasets, DexYCB [5], HO3D [15], H2O3D [16], ARCTIC [12], HOI4D [36], and H2O [24] exhibit significant spatial imbalance, with ground-truth contact concentrated predominantly on the fingertips. ObMan [19] and MOW [4] display more promising contact distributions, including high contact probability on the hypothenar region. However, ObMan is a synthetic dataset with domain gap towards real images, and MOW, while showing diverse palmar contact patterns, lacks dorsal contact and contains fewer than 1,000 samples, making it insufficient for large-scale training. Among hand-hand interaction datasets, HIC [59] shows severe spatial imbalance, with contact limited to small fingertip regions and regions between fingers. In contrast, InterHand2.6M [41] offers a more favorable contact distribution, covering nearly all hand regions except for a small area on the dorsum. Hand-scene interaction datasets [17, 22] tend to involve large contact areas, primarily on the palmar side. However, interactions are often limited to flat surfaces such as ground or walls, resulting in low diversity of contact types despite broad coverage. The hand-face interaction dataset, Decaf [55], also exhibits poor spatial diversity due to its limited action types, such as poking or touching the chin. For the hand-body interaction dataset, Hi4D [65] includes extensive contact regions, but most interactions involve full palmar contact during actions such as hugging or patting. As visualized in the dataset-wide mean contact heatmaps, spatial imbalance is a prevalent issue across existing dense hand contact datasets. To address this, we propose the vertex-level class-balanced (VCB) loss, which mitigates spatial imbalance by reweighting the loss at each vertex based on its mean contact frequency across the dataset. Please refer to the main paper for further details on the VCB loss.
@@
Figure A2: Dataset-wise dense hand contact mean. These heatmaps show average hand contact of ground-truths from ObMan [19], DexYCB [5], MOW [4], HO3D [15], H2O3D [16], ARCTIC [12], HOI4D [36] dataset.
@@
Figure A3: Dataset-wise dense hand contact mean. These heatmaps show average hand contact of ground-truths from H2O [24], InterHand2.6M [41], HIC [59], PROX [17], RICH [22], Decaf [55], Hi4D [65] dataset.
Table A2: Comparison with SOTA methods of dense hand contact estimation on HIC [59], RICH [22], Hi4D [65] dataset.
Test dataset Methods $\text{Precision}{\uparrow}$ $\text{Recall}{\uparrow}$ $\text{F1-Score}{\uparrow}$
HIC [59] POSA [18] 0.000 0.000 0.000
BSTRO [22] 0.000 0.000 0.000
DECO [58] 0.005 0.037 0.006
HACO (Ours) 0.216 0.409 0.263
RICH [22] POSA [18] 0.143 0.175 0.125
BSTRO [22] 0.437 0.498 0.455
DECO [58] 0.324 0.351 0.303
HACO (Ours) 0.741 0.899 0.781
Hi4D [65] POSA [18] 0.128 0.109 0.111
BSTRO [22] 0.313 0.207 0.247
DECO [58] 0.197 0.162 0.146
HACO (Ours) 0.555 0.636 0.565
A.6 More comparisons with state-of-the-art methods
Dense hand contact estimation models. In our main manuscript and Table A2, we provided comparison between our HACO and state-of-the-art (SOTA) methods [18, 22, 58] without any modification on model and its model weights. Nevertheless, the scale of training dataset is largely different between HACO and the SOTA methods. Also, all of the prior methods are dense human-scene contact estimation methods based on human body models such as SMPL [37] and SMPL-X [48]. To provide the comparison that excludes the effect of large-scale training and discrepancy of performance rooted by different human models (e.g., SMPL, SMPL-X, MANO), we further provide comparison with SOTA methods that are trained on the same 14 datasets like HACO and modified as hand version in Table A3. To modify the methods to hand version, we replaced all components of SMPL or SMPL-X modules within POSA [18], BSTRO [22], DECO [58] into MANO [54] modules and changed corresponding model configurations such as the number of vertices accordingly. For POSA, as the method requires estimated 3D mesh, we provide the estimated 3D hand mesh with HaMeR [49]. According to Table A3, HACO still consistently outperforms all hand-specific variants of the full-body methods. These results highlight the strength of HACO’s task-specific design based on balanced contact sampling and vertex-level class-balanced loss.
Table A3: Comparison with SOTA methods of dense hand contact estimation with hand version of POSA [18], BSTRO [22], DECO [58] on MOW [4], HIC [59], RICH [22], Hi4D [65] dataset.
Test dataset Methods $\text{Precision}{\uparrow}$ $\text{Recall}{\uparrow}$ $\text{F1-Score}{\uparrow}$
MOW [4] POSA [18] (hand ver. w/ HaMeR [49] pred. mesh) 0.504 0.400 0.408
BSTRO [22] (hand ver.) 0.518 0.359 0.380
DECO [58] (hand ver.) 0.495 0.385 0.395
HACO (Ours) 0.525 0.607 0.522
HIC [59] POSA [18] (hand ver. w/ HaMeR [49] pred. mesh) 0.060 0.225 0.086
BSTRO [22] (hand ver.) 0.070 0.293 0.107
DECO [58] (hand ver.) 0.067 0.210 0.094
HACO (Ours) 0.216 0.409 0.263
RICH [22] POSA [18] (hand ver. w/ HaMeR [49] pred. mesh) 0.538 0.355 0.398
BSTRO [22] (hand ver.) 0.581 0.413 0.421
DECO [58] (hand ver.) 0.519 0.321 0.367
HACO (Ours) 0.741 0.899 0.781
Hi4D [65] POSA [18] (hand ver. w/ HaMeR [49] pred. mesh) 0.537 0.257 0.315
BSTRO [22] (hand ver.) 0.594 0.375 0.420
DECO [58] (hand ver.) 0.553 0.244 0.304
HACO (Ours) 0.555 0.636 0.565
Dense hand contact estimation modules. While not designed as a standalone dense hand contact estimation model, there are several works that present dense hand contact estimation module as part of 3D reconstruction model [55, 21, 61, 35] or pseudo ground-truth generation of dense hand contact [50]. For contact modules in 3D reconstruction model, they can be divided into learning-based modules [55, 21, 61] and heuristic-driven techniques [35, 50]. These methods were excluded from our SOTA comparison in the main manuscript due to the lack of official code release [55, 21, 61], the task-specific scope of their training data [55, 61], and the heuristic nature of their modules [35, 50]. However, we include a broader evaluation in Table A4 to provide a more comprehensive perspective for future research and to facilitate extensive assessment of hand-specific methods. We first reproduce DefConNet from Decaf [55], Stage 1 of CHOI [21], and InteractionNet from DICE [61] based on the model specifications described in their manuscripts, as neither the code nor the pretrained weights are publicly available. For the Data Engine in LatentAct [50] and EasyHOI [35], we use the officially released code to generate outputs for evaluation. For all learning-based modules [55, 21, 61], we train the reproduced modules based on 14 datasets like HACO but follow their training configurations such as loss function or optimizers when provided in their manuscripts. In Table A4, the results show that our HACO outperforms all previous dense hand contact estimation modules by a significant margin. This highlights the need for a standalone dense hand contact estimation model specifically trained for hand contact, and demonstrates the effectiveness of HACO’s task-specific design.
Table A4: Comparison with SOTA methods of dense hand contact estimation with hand contact estimation modules (†: reproduced, re-trained with 14 datasets, code not available) on MOW [4] dataset.
Methods $\text{Precision}{\uparrow}$ $\text{Recall}{\uparrow}$ $\text{F1-Score}{\uparrow}$
†DefConNet (from Decaf [55]) 0.414 0.404 0.368
†Stage 1 in CHOI [21] 0.521 0.387 0.407
†InteractionNet (from DICE [61]) 0.459 0.392 0.413
Data Engine in LatentAct [50] (w/ SAM2 [51] & HaMeR [49]) 0.345 0.244 0.211
EasyHOI [35] 0.480 0.228 0.282
HACO (Ours) 0.525 0.607 0.522
A.7 More ablation studies
Effectiveness of training dataset size. To assess the contribution of large-scale training to HACO’s performance, we conducted an ablation study comparing different training dataset sizes spanning 1, 3, and 14 datasets. The results demonstrate that performance improves notably as the training dataset size increases, emphasizing the importance of large-scale and diverse data for robust contact estimation. In particular, training on the full set of 14 datasets yields the best performance, while using only three representative datasets of DexYCB [5], ObMan [19], MOW [4] and even one dataset of MOW dataset still produces competitive results, indicating that HACO remains effective even with limited training dataset size.
Table A5: Ablation for training dataset size on MOW [4] dataset.
Methods $\text{Precision}{\uparrow}$ $\text{Recall}{\uparrow}$ $\text{F1-Score}{\uparrow}$
HACO trained on 1 dataset 0.498 0.348 0.373
HACO trained on 3 datasets 0.463 0.602 0.485
HACO trained on 14 datasets 0.525 0.607 0.522
A.8 Quantitative results with different backbones
@@
We evaluate the performance of HACO with various backbone architectures, keeping the rest of the pipeline fixed. As shown in Table A6, our original model using ViT-H [10] pretrained on HaMeR [49] achieves the highest F1-score, demonstrating the strongest overall performance. ViT-based backbones consistently outperform convolutional alternatives, with ViT-B achieving the second highest recall and ViT-S maintaining strong precision. ViT-L also performs competitively, achieving a strong F1-score. This highlights the effectiveness of Transformer [60]-based architectures for dense contact estimation, likely due to their global receptive fields and capacity to model long-range dependencies. To assess the benefit of pretraining on related tasks, we additionally train and evaluate a variant with FPN [31] backbone pretrained with HandOccNet [46], a 3D hand mesh reconstruction model. This configuration yields low F1-score compared to ImageNet-pretrained backbones, suggesting that reconstruction-specific pretraining alone is insufficient for accurate hand contact prediction. HRNet backbones, widely used in 3D human mesh reconstruction [29, 30, 28, 66, 27], achieve moderate performance in our setting, with F1-scores below 0.5 despite their high-resolution design. ResNet backbones [20] exhibit a wide performance range, with deeper variants such as ResNet-152 performing relatively well, while lighter variants such as ResNet-18 offer improved efficiency at the cost of reduced F1-score. Overall, these results indicate that the performance gains of our final HACO model arise from the combination of the strong representational capacity of the ViT backbone and task-specific knowledge transferred from 3D hand mesh reconstruction.
Table A6: Comparison of various backbone models on MOW [4] dataset.
Model Backbone Pretrained $\text{Precision}{\uparrow}$ $\text{Recall}{\uparrow}$ $\text{F1-Score}{\uparrow}$
HACO ViT-H [10] HaMeR [49] 0.525 0.607 0.522
HACO FPN [31] HandOccNet [46] 0.505 0.575 0.482
HACO ViT-L [10] ImageNet [8] 0.493 0.598 0.491
HACO ViT-B [10] ImageNet [8] 0.488 0.604 0.484
HACO ViT-S [10] ImageNet [8] 0.510 0.514 0.462
HACO HRNet-W48 [57] ImageNet [8] 0.509 0.563 0.485
HACO HRNet-W32 [57] ImageNet [8] 0.500 0.558 0.471
HACO ResNet-152 [20] ImageNet [8] 0.506 0.558 0.486
HACO ResNet-101 [20] ImageNet [8] 0.498 0.564 0.480
HACO ResNet-50 [20] ImageNet [8] 0.490 0.558 0.470
HACO ResNet-34 [20] ImageNet [8] 0.489 0.558 0.466
HACO ResNet-18 [20] ImageNet [8] 0.484 0.538 0.456
A.9 Computational requirements